12. A 200 mm focal length lens can focus on objects if they are between 2m and infinity away. What is the range of distances the lens must be from the film/detector? [.2 to .222 m]
Answer #1

Use the lens formula,

1/f = 1/v + 1/u

for u = infinity,

\frac{1}{0.2} = \frac{1}{v}+\frac{1}{\infty}

=> v = 0.2 m = 20 cm

for u = 2 m,

\frac{1}{0.2} = \frac{1}{v}+\frac{1}{2}

=> v = 0.2222 m = 22.22 cm.

So, the range of distance where the lens must be from the detector will be from 0.2 meters to 0.2222 meters.